OpenStudy (anonymous):

What value of k solves the equation? k^-2 = 1/64 A. 1/8 B.8 C. 32 D. 128

OpenStudy (anonymous):

Plug in the answers and see if they work.

OpenStudy (campbell_st):

so when not written in index form the problem is \[\frac{1}{k^2} = \frac{1}{64}\] this should help you find k, and then find the answer

OpenStudy (anonymous):

<3

OpenStudy (jdoe0001):

keep in mind that \(\large \bf a^{-{\color{red} n}} = \cfrac{1}{a^{\color{red} n}}\)

OpenStudy (anonymous):

I think 1/8 or 8..

OpenStudy (campbell_st):

well k^2 = 64 so what is k

OpenStudy (jdoe0001):

\(\bf \large { a^{-{\color{red} n}} = \cfrac{1}{a^{\color{red} n}} \\ \quad \\ \quad \\ k^{-{\color{red}{ 2}}}=\cfrac{1}{64}\implies \cfrac{1}{k^{\color{red}{ 2}}}=\cfrac{1}{64}\implies \begin{array}{llll} k^{\color{red}{ 2}}=64 \\ \quad \\ k=? \end{array} }\)

OpenStudy (anonymous):

8

OpenStudy (jdoe0001):

yeap